# Anomaly induced transport from symmetry breaking in holography

### v1

@@ 6.4.1 Fixed non-vanishing chemical potentials
We will present below the results corresponding to fixing the chemical potentials as $\mu_{v}/(\pi T)=0.1$, $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$. We have computed the conductivities involving vector, axial, non-anomalous and energy currents as a function of the dimensionless mass parameter $M/(\pi T)$. We show the plots of the conductivities normalized by their $M\to 0$ values when these are finite. In cases where the conductivities vanish in this limit, we normalized them by using some related non-vanishing conductivities from the same sector. The results are given in figs. 17,  18 and 19. We can see in these figures that the conductivities inducing the axial current $\vec{J}_{a}$ are monotonically suppressed, while the conductivities inducing the non-anomalous current $\vec{J}_{w}$ become enhanced when $M/(\pi T)$ increases. This indicates a redistribution of the magnetic response from the axial to the non-anomalous sector, while the vector sector is less affected. In addition, all the chiral magnetic conductivities that control the response to a magnetic field in the non-anomalous sector ($\sigma^{B}_{vw}$, $\sigma^{B}_{aw}$, and $\sigma^{B}_{ww}$) increase with $M/(\pi T)$. The chiral vortical conductivities exhibit a complementary structure: the vector part ($\sigma^{V}_{v}$) is less affected by the change in $M/(\pi T)$, the axial part ($\sigma^{V}_{a}$) is suppressed, and the non-anomalous sector ($\sigma^{V}_{w}$) is amplified.
@@ 6.4.3 Results in the large $M$ limit
While both symmetries $A_{+}$ and $A_{-}$ have CS interactions, the scalar field breaks only the $A_{-}$ symmetry. At $M=0$ there will be chiral conductivities for both $A_{\pm}$, but for large $M$, $A_{-}$ will be badly broken, and the chiral magnetic effect in the $J_{-}$ current goes to zero.  [^6] See Ref. Jimenez-Alba:2014iia for a similar effect. This implies the vanishing of $\sigma^{B}_{-v}$ in the limit $M\to\infty$, so that one has from eq. (48) the following relation
@@
The expressions for the other conductivities in eq. (50) can be read off from eqs. (57)-(65). The results for all the conductivities in the large $M$ limit and in the basis $(V,A,W)$ are collected in appendix B. Notice that the values of the conductivities in this limit are not directly related in general to their values in the massless case due to their dependence on $\mu_{+}\equiv\mu_{a}+\mu_{w}$, except for the case studied in eq. (66). In addition, it turns out that in general all the conductivities have different values at $M=0$ and at $M\to\infty$, thus implying that all of them have a dependence on the symmetry breaking parameter $M$. However, there is a single particular non-trivial situation in which some of the conductivities have the same values in both limits, i.e. $\sigma^{B}_{vv}$, $\sigma^{B}_{\varepsilon v}$, $\sigma^{V}_{v}$ and $\sigma^{V}_{\varepsilon}$ for $\mu_{a}=\mu_{w}$ [^7] The trivial situation in which this happens corresponds to $T=0$ and/or $\mu_{s}=0\;\;(s\in\{v,a,w\})$. We have checked that if $\mu_{a}=\mu_{w}$, this is not just an asymptotic property but in fact these four conductivities turn out to be independent of the mass parameter $M$.
We see in figs. 17, 18, 19 and 20 that the tendency of the numerical results is to confirm the relations of eq. (50). A more robust check demands computing the conductivities for higher values of $M/(\pi T)$. However, this problem turns out to be difficult when considering the full backreaction of the metric due to the great increase of the numerical error for larger values of $M$. A way of overcoming this difficulty is to work in the probe limit, as we expect that the computation with no backreaction leads to a good approximation in the limit $M/(\pi T)\gg 1$ for conductivities whose results just depend on the chemical potentials and not on the temperature: this is the case of conductivities related to charged transport induced by magnetic fields, but not those related to the vortical effect or the energy transport sector which demand a computation with full backreaction of the metric. This is confirmed by the numerics. We display in figs. 21 and 22 the numerical results in the probe limit for the conductivities in the charged sector in the range $0\leq M/(\pi T)\leq 15$, together with the corresponding analytical results in the large $M$ limit (cf. appendix B). Note that the agreement between numerics and analytical formulas is excellent for $M/(\pi T)\gtrsim 10$.
@@ 7 Discussion and outlook
In summary, our results reveal how symmetry breaking terms interact explicitly with anomaly contributions to modify the response to anomalous transport. One of the the most interesting results of this work is that once the symmetry is broken by the scalar field $\phi$, even the non-anomalous current $\vec{J}_{w}$ shows response akin to the chiral vortical and chiral magnetic effect. It would be interesting to study the analogous phenomenon in a weak coupling field theoretical setup. One might also think of possible application in condensed matter physics. Weyl semi-metals typically come with a multiple of Weyl nodes (valleys) in the Brillouin zone Basar:2013iaa; Landsteiner:2015lsa. One can speculate that they feature not only axial symmetries by also some non-anomalous accidental valley symmetries and corresponding valley currents. It would be fascinating to see a similar non-anomalous CME in such valley currents in certain phases in which parts of valley symmetries are broken.
@@ Acknowledgements.
A.T. would like to thank his supervisor Shubhrangshu Ghosh for his valuable suggestions and help throughout the work. A.T. acknowledges financial support by the Ministry of Tribal Affairs, Government of India, through the National Fellowship for Scheduled Tribes (NFST). The work of K.L. is partially supported by CEX2020-001007-S and PID2021-123017NB-100, PID2021-127726NB-I00 funded by MCIN/AEI/10.13039/501100011033 and ERDF “A way of making Europe”. The work of E.M. is supported by the “Proyectos de Investigación Precompetitivos” Program of the Plan Propio de Investigación of the University of Granada. N.R. and E.M. acknowledge financial support during the initial stages of this work by the Ramón y Cajal Program under Grant RYC-2016-20678 funded by MCIN/AEI/10.13039/501100011033 and by “FSE Investing in your future”, and by the FEDER/Junta de Andalucía-Consejería de Economía y Conocimiento 2014–2020 Operational Program under Grant A-FQM-178-UGR18.
@@ A.2 Equations of motion of the fluctuations
Appendix B Conductivities in the large $M$ limit

### v2

@@ 6.4.1 Fixed non-vanishing chemical potentials
We will present below the results corresponding to fixing the chemical potentials as $\mu_{v}/(\pi T)=0.1$, $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$. We have computed the conductivities involving vector, axial, non-anomalous and energy currents as a function of the dimensionless mass parameter $M/(\pi T)$. We show the plots of the conductivities normalized by their $M\to 0$ values when these are finite. In cases where the conductivities vanish in this limit, we normalized them by using some related non-vanishing conductivities from the same sector. The results are given in figs. 6,  7 and 8. We can see in these figures that the conductivities inducing the axial current $\vec{J}_{a}$ are monotonically suppressed, while the conductivities inducing the non-anomalous current $\vec{J}_{w}$ become enhanced when $M/(\pi T)$ increases. This indicates a redistribution of the magnetic response from the axial to the non-anomalous sector, while the vector sector is less affected. In addition, all the chiral magnetic conductivities that control the response to a magnetic field in the non-anomalous sector ($\sigma^{B}_{vw}$, $\sigma^{B}_{aw}$, and $\sigma^{B}_{ww}$) increase with $M/(\pi T)$. The chiral vortical conductivities exhibit a complementary structure: the vector part ($\sigma^{V}_{v}$) is less affected by the change in $M/(\pi T)$, the axial part ($\sigma^{V}_{a}$) is suppressed, and the non-anomalous sector ($\sigma^{V}_{w}$) is amplified.
@@ 6.4.3 Results in the large $M$ limit
While both symmetries $A_{+}$ and $A_{-}$ have CS interactions, the scalar field breaks only the $A_{-}$ symmetry. At $M=0$ there will be chiral conductivities for both $A_{\pm}$, but for large $M$, $A_{-}$ will be badly broken, and the chiral magnetic effect in the $J_{-}$ current goes to zero.  [^6] See Ref. Jimenez-Alba et al. (2014) for a similar effect. This implies the vanishing of $\sigma^{B}_{-v}$ in the limit $M\to\infty$, so that one has from eq. (49) the following relation
@@
The expressions for the other conductivities in eq. (51) can be read off from eqs. (58)-(66). The results for all the conductivities in the large $M$ limit and in the basis $(V,A,W)$ are collected in appendix C. Notice that the values of the conductivities in this limit are not directly related in general to their values in the massless case due to their dependence on $\mu_{+}\equiv\mu_{a}+\mu_{w}$, except for the case presented in eq. (67). In addition, it turns out that in general all the conductivities have different values at $M=0$ and at $M\to\infty$, thus implying that all of them have a dependence on the symmetry breaking parameter $M$. However, there is a single particular non-trivial situation in which some of the conductivities have the same values in both limits, i.e. $\sigma^{B}_{vv}$, $\sigma^{B}_{\varepsilon v}$, $\sigma^{V}_{v}$ and $\sigma^{V}_{\varepsilon}$ for $\mu_{a}=\mu_{w}$ [^7] The trivial situation in which this happens corresponds to $T=0$ and/or $\mu_{s}=0\;\;(s\in\{v,a,w\})$. We have checked that if $\mu_{a}=\mu_{w}$, this is not just an asymptotic property but in fact these four conductivities turn out to be independent of the mass parameter $M$.
We see in figs. 6, 7, 8 and 9 that the tendency of the numerical results is to confirm the relations of eq. (51). A more robust check demands computing the conductivities for higher values of $M/(\pi T)$. However, this problem turns out to be difficult when considering the full backreaction of the metric due to the great increase of the numerical error for larger values of $M$. A way of overcoming this difficulty is to work in the probe limit, as we expect that the computation with no backreaction leads to a good approximation in the limit $M/(\pi T)\gg 1$ for conductivities whose results just depend on the chemical potentials and not on the temperature, a property that is confirmed by the numerics: this is the case of conductivities related to charged transport induced by magnetic fields, but not those related to the vortical effect or the energy transport sector which demand a computation with full backreaction of the metric. We display in figs. 10 and 11 the numerical results in the probe limit for the conductivities in the charged sector in the range $0\leq M/(\pi T)\leq 15$, together with the corresponding analytical results in the large $M$ limit (cf. appendix C). Note that the agreement between numerics and analytical formulas is excellent for $M/(\pi T)\gtrsim 10$.
@@ 7 Discussion and outlook
In summary, our results reveal how symmetry breaking terms interact explicitly with anomaly contributions to modify the response to anomalous transport. One of the most interesting results of this work is that once the symmetry is broken by the scalar field $\phi$, even the non-anomalous current $\vec{J}_{w}$ shows response akin to the chiral vortical and chiral magnetic effect. It would be interesting to study the analogous phenomenon in a weak coupling field theoretical setup. One might also think of possible application in condensed matter physics. Weyl semi-metals typically come with a multiple of Weyl nodes (valleys) in the Brillouin zone Basar et al. (2014); Landsteiner and Liu (2016). One can speculate that they feature not only axial symmetries by also some non-anomalous accidental valley symmetries and corresponding valley currents. It would be fascinating to see a similar non-anomalous CME in such valley currents in certain phases in which parts of valley symmetries are broken.
@@ Acknowledgements.
E.M. thanks Grigory E. Volovik for useful correspondence. A.T. would like to thank his supervisor Shubhrangshu Ghosh for his valuable suggestions and help throughout the work. A.T. acknowledges financial support by the Ministry of Tribal Affairs, Government of India, through the National Fellowship for Scheduled Tribes (NFST). The works of N.R. and E.M. are supported by the “Proyectos de Investigación Precompetitivos” Program of the Plan Propio de Investigación of the University of Granada under grant PP2025PP-18. N.R. and E.M. acknowledge financial support during the initial stages of this work by the Ramón y Cajal Program under Grant RYC-2016-20678 funded by MCIN/AEI/10.13039/501100011033 and by “FSE Investing in your future”, and by the FEDER/Junta de Andalucía-Consejería de Economía y Conocimiento 2014–2020 Operational Program under Grant A-FQM-178-UGR18. The work of K.L. is partially supported by CEX2020-001007-S and PID2021-123017NB-I00, PID2021-127726NB-I00, PID2024-156043NB-I00 funded by MCIN/AEI/10.13039/501100011033 and ERDF “A way of making Europe”.
@@ A.2 Equations of motion of the fluctuations
Appendix B Conductivities at finite $M$
We present in this appendix all the numerical results for the conductivities appearing in eqs. (25)-(28) as a function of $\mu_{v}/(\pi T)$, $\mu_{a}/(\pi T)$ and $\mu_{w}/(\pi T)$, for several values of $M/(\pi T)$. When presenting the results, we will classify them in four categories corresponding to the conductivities controlling the response to: i) $\vec{B}_{v}$, ii) $\vec{B}_{a}$, iii) $\vec{B}_{w}$, and iv) $\vec{\Omega}$; in this order. Unless otherwise stated, we will use as fixed values of the chemical potentials the following ones: $\mu_{v}/(\pi T)=0.1$, $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$; and we will consider three different cases: Case 1: changing $\mu_{v}/(\pi T)$, Case 2: changing $\mu_{a}/(\pi T)$, and Case 3: changing $\mu_{w}/(\pi T)$. We will show the results obtained by using the numerical method of sec. 6.1 in the range $0\leq M/(\pi T)\leq 2$. For comparison, we will also display the results obtained in the large $M$ limit by using the analytical formulas that will be derived in sec. 6.4.3 and appendix C.
Figure 12: Case 1: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
B.1 Conductivities controlling the response to $\vec{B}_{v}$
  • Case 1: The results are displayed in fig. 12. We can see that $\sigma_{av}^{B}$, $\sigma_{wv}^{B}$ and $\sigma_{\varepsilon v}^{B}$ grow linearly with $\mu_{v}/(\pi T)$. With the increase $M$ the slopes of $\sigma_{av}^{B}$ and $\sigma_{\varepsilon v}^{B}$ decrease, while the slope of $\sigma_{wv}^{B}$ increases. The conductivity $\sigma_{vv}^{B}$ turns out to be independent of $\mu_{v}/(\pi T)$ independently of the value of $M$, exactly as in the absence of symmetry breaking, cf. eq. (37) Landsteiner et al. (2014).
    Figure 13: Case 2: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 14: Case 3: Plots of the conductivities $\sigma^{B}_{sv}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
B.2 Conductivities controlling the response to $\vec{B}_{a}$
Figure 15: Case 1: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
Figure 16: Case 2: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 17: Case 3: Plots of the conductivities $\sigma^{B}_{sa}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
B.3 Conductivities controlling the response to $\vec{B}_{w}$
  • Case 2: The results are displayed in fig. 19. It can be seen that $\sigma^{B}_{vw}$ remains constant with $\mu_{a}/(\pi T)$, but its value increases steadily with $M$. The conductivities $\sigma^{B}_{aw}$ and $\sigma^{B}_{ww}$ grow linearly with $\mu_{a}/(\pi T)$, with their slopes enhanced as $M$ increases, except for the case $\sigma^{B}_{aw}$ and very large values of $M$, as it was discussed in Case 2 of sec. B.2. Finally, $\sigma^{B}_{\varepsilon w}$ shows a quadratic dependence on $\mu_{a}/(\pi T)$, with an increasing curvature as $M$ increases, similar to Case 1.
  • Case 3: The results are displayed in fig. 20. The conductivity $\sigma^{B}_{vw}$ remains independent of $\mu_{w}/(\pi T)$, but it value increases steadily with $M$. The conductivities $\sigma^{B}_{aw}$ and $\sigma^{B}_{ww}$, on the other hand, grow linearly with $\mu_{w}/(\pi T)$, with their slopes enhanced as $M$ increases. The intersection observed for $\sigma^{B}_{aw}$ in the top-right panel of fig. 20 has the same discussion as in Case 3 of sec. B.2. Finally, $\sigma^{B}_{\varepsilon w}$ has a linear ++ quadratic dependence on $\mu_{w}/(\pi T)$ with higher curvature for higher values of $M$. For very large values of $M$, this dependence is of the form $\sigma^{B}_{\varepsilon w}\propto(\mu_{a}+\mu_{w})^{2}$, as it can be seen in eq. (97).
Figure 18: Case 1: Plots of the conductivities $\sigma^{B}_{sw}$, where $S\in\{v,a,w,\varepsilon\}$, as a function of to $\mu_{v}/(\pi T)$. We have set $\mu_{a}/(\pi T)=0.3$ and $\mu_{w}/(\pi T)=0.2$.
Figure 19: Case 2: Plots of the conductivities $\sigma^{B}_{sw}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of to $\mu_{a}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{w}/(\pi T)=0.2$.
Figure 20: Case 3: Plots of the conductivities $\sigma^{B}_{sw}$, where $s\in\{v,a,w,\varepsilon\}$, as a function of $\mu_{w}/(\pi T)$. We have set $\mu_{v}/(\pi T)=0.1$ and $\mu_{a}/(\pi T)=0.3$.
B.4 Conductivities controlling the response to $\vec{\Omega}$
Appendix C Conductivities in the large $M$ limit